A consumer buys an Iphone for Php12,000 and gets consumer surplus of Php8,000.

a. What is the consumer surplus?

b. If the consumer had bought the Iphone on sale for Php9000, what would the consumer surplus have been?

c. If the price of an Iphone were Php2500, what would the consumer surplus have been?


1
Expert's answer
2021-10-28T17:38:27-0400

a)

CS1=12000-8000=4000 Php,

b)

CS2=9000-8000=1000 Php,

c)

CS3=12000-2500=9500 Php.